Session: Sep 1998 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M.Rioja at Thu Sep 10 09:46:58 METDST 1998) Minor failures
Severe problems at the beginning. First succesful observations at 23:30 UT - lost 20 minutes. Weather: Clear sky.

Onsala (A. Polatidis at Thu Sep 10 10:44:04 METDST 1998) Success
Tape lost vacuum at 2314, repositioned and started tape at 2318. No other problems with the data. Very strong winds during most of the experiment (11-13 m/s). Rain for most part of the night.

Westerbork (Tony Foley at Thu Sep 10 15:00:48 METDST 1998) Success
Using RT7 (single dish) and Rubidium clock (maser not working). clock offset about 2.5 microseconds. Data looks reasonable; measured Tsys about 65-70K in a warm (16C) humid night

Noto (G. Tuccari at Tue Sep 15 13:25:47 METDST 1998) Success
No particular problem has been reported by operators. Clear sky.

Urumqi (Liu Xiang at Wed Sep 16 05:12:59 METDST 1998) Minor failures
It was clear, the clock offset was 5.4us. The computer problem in UT: 00:03:00--00:10:00, so four scans lost, and in UT: 03:22:00 ---03:23:04, tape recorder problem. No other problem.

Effelsberg (dgraham at Fri Sep 18 15:58:25 METDST 1998) Minor failures
low elevation until 00 UT, below local horizon

Cambridge (Alastair Gunn at Fri Sep 18 17:39:17 METDST 1998) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Mon Sep 21 17:22:26 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the sep98 directory under the name et004a.antab. Read the comments in the file to get further information.